OpenStudy (anonymous):

What is the simplified form of (x^2+5x+6)/(15xy^2)/(2x^2+7x+3)/(5x^2y) I got x(x-2)/3y(2x-y) for my answer

OpenStudy (anonymous):

@Dhawal

OpenStudy (anonymous):

@phi

OpenStudy (phi):

did you factor the expressions?

OpenStudy (anonymous):

i tried to but i dont know if i did it correctly

OpenStudy (phi):

if you can you should use the equation editor, because these are hard to read after "flipping" and multiplying you have \[ \frac{(x^2+5x+6}{(15xy^2)} \frac{(5x^2y)} {(2x^2+7x+3)}\]

OpenStudy (phi):

notice you can simplify the "easy part" \[ \frac{5x^2y}{15xy^2} = \frac{x}{3y}\]

OpenStudy (phi):

x^2+5x+6 factors into (x+2)(x+3) 2x^2+7x+3 factors into (2x+ 1 )(x+3) the (x+3) will cancel leaving (x+2)/(2x+1) you get as a final answer \[ \frac{x(x+2)}{3y(2x+1)} \] which is not what you posted as an answer